A 2.0 kg mass has 300 J energy. What is the speed of the mass


1
Expert's answer
2020-11-20T06:53:55-0500

The kinetic energy of the body is

"\displaystyle E = \frac{mv^2}{2} = 300 \; J" .

The mass is given: "m=2.0 \; kg".

So, the speed is

"\displaystyle v = \sqrt{\frac{2E}{m}} = \sqrt{\frac{600}{2}} = \sqrt{300} \approx 17.3 \; \frac{m}{s}."

Answer: 17.3 m/s.